# R-Bot: An LLM-based Query Rewrite System

### v1

@@ 7. Experiments
We conduct extensive experiments to evaluate our system R-Bot and compare with state-of-the-art query rewrite systems.
@@ 7.1. Experiment Setting
We implement our system R-Bot using the rules in an open-sourced query engine Apache Calcite (Begoli et al., 2018). We execute the SQL queries in PostgreSQL v14. The server is a machine with 125.6 GB RAM and 3.1GHz CPU.
Datasets. To verify the effectiveness of R-Bot on different scenarios, we conduct experiments on three types of datasets. $(i)$ TPC-H is a standard OLAP benchmark, which contains 62 columns and 44 queries. We separately test R-Bot on different data sizes, i.e., TPC-H 10x ($\sim$10G) and TPC-H 50x ($\sim$50G). $(ii)$ DSB is a more complex OLAP benchmark (Ding et al., 2021), which contains 429 columns and 76 queries. We also test R-Bot on datasets DSB 10x ($\sim$10G) and DSB 50x ($\sim$50G). $(iii)$ Calcite is constructed from Apache Calcite’s rewrite rule test suites (cal, 2024), which is widely used in evaluating query rewrite capabilities (Wang et al., 2022; Dong et al., 2023; Bai et al., 2023). In our experiment, we select 44 representative queries with great rewrite potentials on a schema of 43 columns.  [^1] We chose queries from the Calcite test suites that showed the top 11% biggest performance improvements when rewritten by the Calcite query rewriter. Additionally, we evaluate R-Bot with a 10G data size across distinct data distributions, namely Calcite (uni) for uniform distribution and Calcite (zipf) for Zipfian distribution.
@@
Evaluated Methods. The evaluated methods include: $(i)$ R-Bot (GPT-4) is the version of R-Bot using gpt-4o. $(ii)$ R-Bot (GPT-3.5) is the version of R-Bot using gpt-3.5-turbo-0125. $(iii)$ LearnedRewrite employs Monte Carlo Tree Search (MCTS) to explore the optimal sequence of rewrite rules (Zhou et al., 2021). It estimates the performance improvements of the rewritten queries by utilizing database query costs, which guides the search process.  [^2] The origin LearnedRewrite learns neural networks from historical queries to estimate the cost reduction of applying rewrite rules. However, as our experiments are conducted in a practical setting without training examples, we replace the learning-based cost model with database query costs, which is more accurate while incurring longer rewrite latency. $(iv)$ GPT-4 model without using techniques in R-Bot, which inputs the SQL query and the rewrite rules, and outputs the arranged rule sequence. $(v)$ GPT-3.5 model similarly without using techniques in R-Bot.
@@ 7.2. Performance Comparison
First, R-Bot can judiciously retrieve relevant rule specifications and Q&As, and use them to comprehensively figure out the rewrite rules pertinent to the input query. For instance, R-Bot can utilize the rewrite evidences to decide to apply the exploration rule “EXPAND_DISTINCT_AGGREGATES_TO_JOIN” weakly related to the input query, which results in a query rewrite of 5.6x acceleration. However, the other methods do not figure out this critical rule without guidance of rewrite evidences. Besides, we find that GPT-4 and GPT-3.5 tend to select only a small number of rules (e.g., 1), which is often sub-optimal. That is because, they lack of rewrite evidences, and thus can hardly uncover the intricate correlations between the rewrite rules and the input query.
@@
Figure 5. Comparison of Overall Latency on TPC-H 50x.
Figure 6. Comparison of Overall Latency on DSB 50x.
Overall Latency Reduction. We also evaluate the overall latency of the methods on larger-scale datasets TPC-H 50x and DSB 50x. Figure 5 and Figure 6 demonstrate the results. We find that R-Bot achieves the best overall latency on the two datasets. The reasons are three-fold. First, R-Bot can still optimize the query latency by finding better plans, thus outperforming other methods. Second, R-Bot takes some time (e.g., average around 1 min) to retrieve beneficial evidences and iteratively instruct LLM to select and arrange rewrite rules. Thus it can steadily identify the critical rewrite rules (e.g., “FILTER_SUB_QUERY_TO_JOIN”) with much higher query latency reduction. Instead, other methods do not spend efforts understanding rewrite evidences, which often causes sub-optimal rewritten queries. Third, compared with results of TPC-H 10x and DSB 10x (see Table 3), we find that R-Bot remains similar rewrite latency but demonstrates higher query latency reduction proportional to data scale, which both results in superior overall latency.
@@
LearnedRewrite
6/44
(13.6%)
9/76
(11.8%)
36/44
(81.8%)
GPT-3.5
3/44
(6.8%)
4/76
(5.3%)
16/44
(36.4%)
GPT-4
6/44
(13.6%)
4/76
(5.3%)
21/44
(47.7%)
R-Bot (GPT-3.5)
21/44
(47.7%)
16/76
(21.0%)
31/44
(70.4%)
R-Bot (GPT-4)
17/44
(38.6%)
19/76
(25.0%)
39/44
(88.6%)
@@
Table 5. Comparison of Query Latency on Calcite (zipf).
Query Latency (s) Calcite (zipf)
Average Median p90
Origin 106.31 37.91 300.00
LearnedRewrite
15.94
($\downarrow$85.0%)
4.96
($\downarrow$86.9%)
34.96
($\downarrow$88.3%)
GPT-3.5
58.33
($\downarrow$45.1%)
20.04
($\downarrow$47.1%)
300.00
($\downarrow$0.0%)
GPT-4
61.80
($\uparrow$41.9%)
14.15
($\downarrow$62.7%)
300.00
($\uparrow$0.0%)
R-Bot (GPT-3.5)
32.44
($\downarrow$69.5%)
6.58
($\downarrow$82.6%)
57.40
($\downarrow$80.9%)
R-Bot (GPT-4)
7.56
($\downarrow$92.9%)
4.96
($\downarrow$86.9%)
18.08
($\downarrow$94.0%)
Robustness. Table 5 demonstrates the robustness of query rewrite methods on the Calcite dataset, which transfers from uniform distribution to zipf distribution. We have two observations. First, compared with Calcite (uni) (see Table 3), the performance of R-Bot remains consistent. R-Bot still achieves the lowest average, median, and p90 of query latency among the methods. That is because the rewrite evidences can cover different query rewrite scenarios, and R-Bot can adaptively decide among potential rule arrangements with feedbacks of database query cost. Second, we find that LearnedRewrite achieves comparative median query latency with R-Bot on Calcite (zipf) dataset. That is because the queries in the Calcite dataset often have great rewrite potentials of applying the rules, where LearnedRewrite can easily find beneficial rule orders even with blind explorations.
Evaluation Across Various LLMs. As shown in Table 3, we find that R-Bot (GPT-3.5) achieves comparative performance with R-Bot (GPT-4). That is because the algorithm design of R-Bot always decomposes complex tasks into simpler stages, both in evidence retrieval and step-by-step LLM rewrite, where each stage can be manageable with inferior LLM (e.g., GPT-3.5).
@@ 7.3. Ablation Study
Table 6. Ablation Study of Rewrite Evidence.
Query Latency (s) Calcite (uni)
Average Median p90
@@
Figure 7. Distribution of Rewrite Evidences with Different SQL Operations.
Table 7. Ablation Study of Structure-Semantics Retrieval.
Query Latency (s) Calcite (uni)
Average Median p90
@@
Table 8. Ablation Study of Step-by-Step LLM Rewrite.
Query Latency (s) Calcite (uni)
Average Median p90
@@
7.3.1. Rewrite Evidence Preparation.
Besides, we also investigate the coverage of prepared rewrite evidences in Figure 7, where 67 rewrite rule specifications are extracted (30 from 119 chapters of documents, 37 from rule codes), and 2091 rewrite Q&As are filtered out from millions of website Q&As. We find that the rule specifications and Q&As both involve a wide range SQL operations (e.g., joins, sub-queries), and can thus provide supportive evidence for various query rewrite scenarios.
7.3.2. Hybrid Structure-Semantics Evidence Retrieval.
To evaluate the effectiveness of hybrid structure-semantics Q&A retrieval, we evaluate R-Bot separately with semantics-only and structure-only retrieval. As indicated in Table 7 , both methods exhibit a decline in performance compared to R-Bot. This reduction is due to their inability to identify Q&As beneficial for query rewrite from a holistic perspective, as they either overlook query semantics or structure information. For example, compared with R-Bot, they may miss key evidence, leading to the omission of critical rules like “EXPAND_DISTINCT_AGGREGATES” in the rule arrangement.
Retrieval Top-k. We also evaluate the impact of retrieval top-$k$, which decides the number of retrieved Q&As. As shown in Figure 9, we have two observations with varying values of $k$. First, the rewrite latency increases as $k$ grows. This is because R-Bot tends to generate more recipes with a greater number of retrieved Q&As. Consequently, as the recipes compose longer LLM contexts, each invocation of LLM requires additional time to process the extended context. Second, as $k$ increases, query latency generally decreases between 1 and 10, but increases between 10 and 20. The reasons are two-fold. $(i)$ When $k$ is small, the additional retrieved Q&As are more likely to cover essential evidence, allowing R-Bot to select critical rules (e.g., “AGGREGATE_JOIN_TRANSPOSE”) to facilitate query rewrite. $(ii)$ For large $k$, it is likely to already cover the essential evidence in the retrieved Q&As, making further increase unnecessary as it may not provide additional relevant information. Besides, increasing $k$ too much may introduce noisy contexts, which may impair performance of LLM in query rewrite. Therefore, we set $k=10$ in the other experiments.
@@
We compare our step-by-step LLM rewrite with one-step LLM rewrite that directly instructs LLM to arrange the rules in one step using retrieved rule specifications and Q&As, and evaluate the performance. As shown in Table 8, we find that one-step LLM rewrite often selects a smaller number of rules, which results in sub-optimal rule arrangements. That is because, given the extremely long context of retrieved evidences and available rewrite rules, LLM tends to produce short-cut solutions overlooking many critical details, and thus only selects the rules explicitly related to the input query. Instead, R-Bot employs a step-by-step LLM rewrite for reasoning, enabling the selection of higher-quality rules.
@@ 8. Related Work
Query Rewrite. Existing query rewrite methods mainly adopt two paradigms. (1) Heuristic-based methods (mys, 2024; pos, 2024; Graefe and McKenna, 1993; Graefe, 1995; Begoli et al., 2018). Some methods (e.g., PostgreSQL (pos, 2024)) apply rewrite rules in a fixed order, which often overlook better orders in different scenarios. Besides, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to explore different rule orders with the aid of heuristic acceleration. However, by neglecting inter-dependencies among the rules, these methods perform a blind search, often failing to identify the optimal orders within reasonable time. (2) Learning-based methods. LearnedRewrite (Zhou et al., 2021) decides the rule order using Monte Carlo Tree Search guided by learned cost models. However, the learned model cannot be transferred to unseen database schema without additional retraining, which is often impractical. Hence, it calls for a query rewrite system capable of reliably identifying an appropriate sequence of rewrite rules to optimize query rewrite.
Automatic Rewrite Rule Discovery. Given the tedious effort required to curate a large number of high-quality rewrite rules, some methods have been recently proposed to automate the process of rule discovery (Wang et al., 2022; Dong et al., 2023; Ding et al., 2023). WeTune (Wang et al., 2022) first represents the condition and transformation of rewrite rule with operators (e.g., projection) and symbols (e.g., relation, attribute). Then it enumerates potential rewrite rules within a constrained set of transformations (e.g., involving up to four operators) and conditions (combinations of seven basic types), verifying rule equivalence using SMT-solvers (Zhou et al., 2022; Chu et al., 2017). Besides, SlabCity (Dong et al., 2023) discovers possible query rewrites for a specific query by recursively applying various transformations (e.g., expanding SQL conditions) to the original query, and preserves the equivalent rewrites. Complementary to the above works, R-Bot retrieves highly relevant evidences to instruct LLM, effectively arranges newly discovered rules, and achieves superior adaptability and robustness in rule selection and ordering.
LLM for Database. Recent advances in LLMs have demonstrated superiority in natural language understanding and programming, which offers opportunities to leverage vast amounts of database forums to enhance database performance (Achiam et al., 2023; Liu et al., 2024b; Fan et al., 2024; Lao et al., 2024; Zhou et al., 2024a, b). For instance, GPTuner (Lao et al., 2024) proposes to enhance database knob tuning using LLM by leveraging domain knowledge to identify important knobs and coarsely initialize their values for subsequent refinement. Besides, D-Bot (Zhou et al., 2024a) proposes an LLM-based database diagnosis system, which can retrieve relevant knowledge chunks and tools, and uses them to accurately identify typical root causes. Unlike these approaches, R-Bot is a novel query rewrite system powered by LLM.

### v2

@@ 7.1. Experiment Setting
We implement our system R-Bot using the rules in an open-sourced query engine Apache Calcite (Begoli et al., 2018). We execute SQL queries in PostgreSQL v14 on a machine with 128 GB RAM and 3.1GHz CPU.
Datasets. To verify the effectiveness of R-Bot on different scenarios, we conduct experiments on three types of datasets. $(i)$ TPC-H is a standard OLAP benchmark, which contains 62 columns and 44 queries. We separately test R-Bot on different data sizes, i.e., TPC-H 10x ($\sim$10G) and TPC-H 50x ($\sim$50G). $(ii)$ DSB is a more complex OLAP benchmark adapted from TPC-DS (Ding et al., 2021), which contains 429 columns and 76 queries. We also test R-Bot on datasets DSB 10x ($\sim$10G) and DSB 50x ($\sim$50G). $(iii)$ Calcite is constructed from Apache Calcite’s rewrite rule test suites (cal, 2024), which is widely used in evaluating query rewrite capabilities (Wang et al., 2022; Dong et al., 2023; Bai et al., 2023). In our experiment, we select 44 representative queries with great rewrite potentials on a schema of 43 columns.  [^2] For efficiency, we chose Calcite test queries with top 11% performance gains under Calcite rewrite. Additional experiments on full queries confirmed that R-Bot consistently reduces query latency and outperforms baseline methods. Additionally, we evaluate R-Bot with a 10G data size across distinct data distributions, namely Calcite (uni) for uniform distribution and Calcite (zipf) for Zipfian distribution.
Figure 5. Distribution of Rewrite Evidences.
Figure 6. Comparison of Overall Latency.
Rewrite Evidences. Our prepared rewrite evidences are shown in Figure 5, including 67 expert-verified rewrite rule specifications (30 from 80k+ tokens of documents, 37 from 30k+ lines of code), and 2091 filtered rewrite Q&As from millions of website Q&As. Together, they cover diverse SQL operations (e.g., joins, sub-queries) and support a wide range of rewrite scenarios.
@@
Evaluated Methods. The evaluated methods include: $(i)$ R-Bot (GPT-4) is R-Bot using gpt-4o. $(ii)$ R-Bot (GPT-3.5) is R-Bot using gpt-3.5-turbo-0125. $(iii)$ LearnedRewrite employs Monte Carlo Tree Search (MCTS) to explore the optimal sequence of rewrite rules (Zhou et al., 2021). It estimates the performance improvements of the rewritten queries with query cost models, which guides the search process.   [^3] Learning-based query rewrite methods, including LLM-based approaches that require model training (Li et al., 2024c), are generally unsuitable for our tests due to the absence of a train-test dataset split. As for LearnedRewrite, we use the authoritative version (lrr, 2024), integrating query cost models approximated using query plan statistics. $(iv)$ GPT-4 without using techniques in R-Bot, which inputs the SQL query and the rewrite rules, and outputs the arranged rule sequence. $(v)$ GPT-3.5 similarly without using techniques in R-Bot.
@@ 7.2. Performance Comparison
First, R-Bot can judiciously retrieve relevant rule specifications and Q&As to figure out the rewrite rules pertinent to the input query. For instance, R-Bot can apply the exploration rule “AGGREGATE_EXPAND_DISTINCT_AGGREGATES_TO_JOIN” weakly related to the input query, using retrieved evidences to achieve a 5.6x optimization. However, the other methods fail to identify this critical rule without guidance from rewrite evidences. Besides, we find that GPT-4 and GPT-3.5 tend to select only a small number of rules (e.g., 1), which is often sub-optimal. That is because, they exhibit hallucination without rewrite evidences, and thus miss the intricate correlations between rewrite rules and the input query.
@@
Overall Latency Reduction. We also evaluate overall latency on larger TPC-H 50x and DSB 50x datasets. As shown in Figure 6, R-Bot achieves the best latency, demonstrating improvements of 1.82x and 1.68x respectively. The reasons are three-fold. First, R-Bot can still optimize the query latency by finding better plans, thus outperforming other methods. Second, R-Bot takes some time (e.g., average around 1 min) to retrieve beneficial evidences and iteratively instruct LLM to select and arrange rewrite rules. Thus it can steadily identify the critical rewrite rules (e.g., “FILTER_SUB_QUERY_TO_JOIN”) with much higher query latency reduction. Instead, other methods do not spend efforts understanding rewrite evidences, which often causes sub-optimal rewritten queries. Third, compared with results of TPC-H 10x and DSB 10x (see Table 4), we find that R-Bot remains similar rewrite latency but demonstrates higher query latency reduction proportional to data scale, which both results in superior overall latency.
@@
LearnedRewrite 7/44 (15.9%) 4/76 (5.3%) 29/44 (63.6%)
GPT-3.5 3/44 (6.8%) 4/76 (5.3%) 16/44 (36.4%)
GPT-4 6/44 (13.6%) 4/76 (5.3%) 21/44 (47.7%)
R-Bot (GPT-3.5) 21/44 (47.7%) 16/76 (21.0%) 31/44 (70.4%)
R-Bot (GPT-4) 17/44 (38.6%) 18/76 (23.7%) 39/44 (88.6%)
Table 6. Comparison of Query Latency on Calcite (zipf).
Query Latency (s) Average Median p90
Origin 106.31 37.91 300.00
LearnedRewrite 71.24 ($\downarrow$33.0%) 5.04 ($\downarrow$86.7%) 300.00 ($\downarrow$0.0%)
GPT-3.5 58.33 ($\downarrow$45.1%) 20.04 ($\downarrow$47.1%) 300.00 ($\downarrow$0.0%)
GPT-4 61.80 ($\uparrow$41.9%) 14.15 ($\downarrow$62.7%) 300.00 ($\downarrow$0.0%)
R-Bot (GPT-3.5) 32.44 ($\downarrow$69.5%) 6.58 ($\downarrow$82.6%) 57.40 ($\downarrow$80.9%)
R-Bot (GPT-4) 7.56 ($\downarrow$92.9%) 4.96 ($\downarrow$86.9%) 18.08 ($\downarrow$94.0%)
@@
Zero-Shot Robustness. First, as shown in Table 4, R-Bot outperforms the other query write methods on different datasets without any re-training. This is because the generalizability of pre-trained LLM enables R-Bot to automatically adapt to unseen database schema and workload. Besides, Table 6 further demonstrates the robustness of query rewrite methods on data distribution, which transfers the Calcite dataset from uniform distribution to zipf distribution. Compared with Calcite (uni) (see Table 4), we find that the performance of R-Bot remains consistent. R-Bot still achieves the lowest average, median, and p90 of query latency among the methods. That is because the rewrite evidences can cover different query rewrite scenarios, and R-Bot can adaptively decide among potential rule arrangements with feedbacks of database query cost.
Evaluation Across Various LLMs. As shown in Table 4, on most metrics, R-Bot (GPT-3.5) performs worse than R-Bot (GPT-4), but still outperforms the other query rewrite methods. That is because R-Bot decomposes complex tasks into simpler stages, both in evidence retrieval and step-by-step LLM rewrite, where each stage can be still manageable even with less advanced LLMs like GPT-3.5.
@@ 7.3. Ablation Study
7.3.1. Rewrite Evidence Preparation.
7.3.2. Structure-Semantics Evidence Retrieval.
To evaluate the effectiveness of hybrid structure-semantics Q&A retrieval, we evaluate R-Bot separately with semantics-only and structure-only retrieval. As indicated in Table 8 , both methods exhibit a decline in performance compared to R-Bot. This reduction is due to their inability to identify Q&As beneficial for query rewrite from a holistic perspective, as they either overlook query semantics or structure information. For example, compared with R-Bot, they may miss key evidence, leading to the omission of critical rules like “AGGREGATE_EXPAND_DISTINCT_AGGREGATES” in arrangement.
Table 7. Ablation Study of Rewrite Evidence on Calcite (uni).
Query Latency (s) Average Median p90
@@
Table 8. Ablation Study of Evidence Retrieval on Calcite (uni).
Query Latency (s) Average Median p90
@@
Table 9. Ablation Study of LLM Rewrite on Calcite (uni).
Query Latency (s) Average Median p90
@@
Retrieval Top-k. We also evaluate the impact of retrieval top-$k$, which decides the number of retrieved Q&As. As shown in Figure 8, we have two observations. First, rewrite latency grows with $k$, because R-Bot tends to generate more recipes with a greater number of retrieved Q&As. Consequently, as the recipes compose longer LLM contexts, each invocation of LLM requires additional time to process the extended context. Second, query latency decreases when $k$ increases from 1 to 10, but rises again beyond 10. The reasons are two-fold. $(i)$ When $k$ is small, the additional retrieved Q&As are more likely to cover essential evidence, allowing R-Bot to select critical rules (e.g., “AGGREGATE_JOIN_TRANSPOSE”). $(ii)$ For large $k$, it is likely to already cover the essential evidence in the retrieved Q&As, making further increase unnecessary as it may not provide additional information. Besides, increasing $k$ too much may introduce noisy contexts, which may impair performance of LLM in query rewrite. Therefore, we set $k=10$ for the other experiments. Meanwhile, we achieve a Precision@10 of 19.8%, indicating that on average at least one relevant Q&A is successfully retrieved.
@@
We compare our step-by-step LLM rewrite with one-step LLM rewrite that instructs LLM to arrange the rules in one step using retrieved evidences, and evaluate the performance. As shown in Table 9, we find that one-step LLM rewrite often selects a smaller number of rules, which results in sub-optimal rule arrangements. That is because, given the extremely long context of retrieved evidences and available rewrite rules, LLM tends to produce short-cut solutions overlooking many critical details, and thus only selects the rules explicitly related to the input query. Instead, R-Bot employs a step-by-step LLM rewrite for reasoning, enabling the selection of higher-quality rules.
@@
Figure 9. Example Query Rewrite Results.
7.4. Case Study
We provide some representative examples to illustrate the query rewrite results. As shown in Figure 9, R-Bot outperforms other methods by accurately identifying critical rules “FILTER_INTO_JOIN” and “FILTER_SUB_QUERY_TO_JOIN” in the first case study, which pushes down conditions involving only one join table and removes the time-consuming sub-queries in the query. On the contrary, GPT-4 selects no useful rewrite rules due to absence of rewrite evidences and dedicated LLM algorithm design. Besides, LearnedRewrite tends to select an excessive number of rewrite rules under erroneous guidance, leading to a less efficient rewritten query. In the second case study, R-Bot also demonstrates superior performance by selecting the rewrite rule “JOIN_CONDITION_PUSH”, which derives an additional condition “$e.deptno=e.empno$”. This condition can be pushed down into the sub-query involving the table “$emp$”, a refinement overlooked by other query rewrite methods. However other methods cannot detect this rule.
7.5. Deployment at Huawei with Real Customers
We have also deployed R-Bot at Huawei, which is empowered by open-sourced LLM DeepSeek-R1-Distill-Qwen-32B (dee, 2024) and embedding model gte-Qwen2-1.5B-instruct (gte, 2024). We verify its effectiveness using a real-world dataset on the largest bank in China, containing 190 columns and 74GB data, as well as 20 real-world slow queries with 7 joined tables and 3 sub-queries on average. As shown in Figure 10, we find that R-Bot can effectively optimize these queries, i.e., improving the latency of 70% queries and reducing the overall latency from 9.23 hours to 4.37 hours. Notably, R-Bot remains effective even when $(i)$ the tested queries are absent from the LLM pre-training corpus (not publicly available) and $(ii)$ open-sourced 32B LLM and embedding model are used as the backend.
@@ 8. Related Work
Query Rewrite. Existing query rewrite methods mainly adopt two paradigms. (1) Heuristic-based methods (mys, 2024; pos, 2024; Graefe and McKenna, 1993; Graefe, 1995; Begoli et al., 2018). Some methods (e.g., PostgreSQL (pos, 2024)) apply rewrite rules in a fixed order, which often overlook better orders in different scenarios. Besides, other heuristic-based methods (e.g., Volcano (Graefe and McKenna, 1993)) attempt to explore different rule orders with the aid of heuristic acceleration. However, by neglecting inter-dependencies among the rules, these methods perform a blind search, often failing to identify the optimal orders within reasonable time. (2) Learning-based methods (Zhou et al., 2021; Li et al., 2024c). LearnedRewrite decides the rule order using Monte Carlo Tree Search guided by learned cost models. However, the learned model cannot be transferred to unseen database schema without additional retraining, which is often impractical. Hence, it calls for a query rewrite system capable of reliably identifying an appropriate sequence of rewrite rules to optimize query rewrite. A recent LLM-based query rewrite method addresses hallucination by incorporating query rewrite examples into the context for LLM to emulate (Li et al., 2024c). However, it still lacks robustness, as it depends on an example pool generated from the training dataset and utilizes a trained model to select examples based on SQL query similarity.
Figure 10. Performance of R-Bot on Real Dataset.
Automatic Rewrite Rule Discovery. Given the tedious effort required to curate a large number of high-quality rewrite rules, some methods have been recently proposed to automate the process of rule discovery (Wang et al., 2022; Dong et al., 2023; Ding et al., 2023). WeTune (Wang et al., 2022) first represents the condition and transformation of rewrite rule with operators and symbols. Then it enumerates potential rewrite rules and verifies rule equivalence using SMT-solvers (Zhou et al., 2022; Chu et al., 2017). SlabCity (Dong et al., 2023) discovers possible query rewrites for a specific query by recursively applying various transformations to the original query. Complementary to the above works, R-Bot retrieves highly relevant evidences to instruct LLM, effectively arranges newly discovered rules, and achieves superior adaptability and robustness in rule selection and ordering.
LLM for Database. There are also many studies to use LLMs to optimize databases (Achiam et al., 2023; Liu et al., 2024b; Fan et al., 2024; Lao et al., 2024; Zhou et al., 2024a; Sun et al., 2025b; Zhou et al., 2024b; Liu and Mozafari, 2024; Zhou et al., 2025c). For instance, GPTuner (Lao et al., 2024) enhances database knob tuning using LLM by leveraging domain knowledge to identify and tune knobs. D-Bot (Zhou et al., 2024a; Sun et al., 2025b) is an LLM-based database diagnosis system. Unlike these approaches, R-Bot is a novel query rewrite system powered by LLM.